Mr. Cantoni invested 50,000.00. A part of it is invested in a bank at 2% yearly interest and another part of it in a mutual fund at a 5% yearly interest. How much investment was made in the mutual fund if his yearly income from the two is 2,800.00?
Found 2 solutions by mananth, MathTherapy:
Answer by mananth(16946)


Part I 2.00% per annum ------------- Amount invested =x
Mutual Fund 5.00% per annum ------------ Amount invested = y =
50000
Interest----- 2800

Part I 2.00% per annum ---x
Mutual Fund 5.00% per annum ---y
Total investment
x + 1 y= 50000 -------------1
Interest on both investments
2.00% x + 5.00% y= 2800
Multiply by 100
2 x + 5 y= 280000.00 --------2
Multiply (1) by -2
we get
-2 x -2 y= -100000.00
Add this to (2)
0 x 3 y= 180000
divide by 3
y = 60000
Part I 2.00% $ -10000
Mutual Fund 5.00% $ 60000

CHECK
-10000 --------- 2.00% ------- -200.00
60000 ------------- 5.00% ------- 3000.00
Total -------------------- 2800.00

m.ananth@hotmail.ca

Answer by MathTherapy(10549)


Mr. Cantoni invested 50,000.00. A part of it is invested in a bank at 2% yearly interest and another part of it in a mutual fund at a 5% yearly interest. How much investment was made in the mutual fund if his yearly income from the two is 2,800.00?
At these 2 rates? IMPOSSIBLE!